# Learning to Ask: When LLMs Meet Unclear Instruction

### v2

@@ Introduction
AwN is built upon widely-used tool-using methods, such as CoT and ReAct. As illustrated in Figure  2, we introduce an additional step before the generation of API calls. This step involves presenting all available information to the LLMs, including the user query and API guideline, and prompting them to determine the adequacy and correctness of user instruction. If LLMs identify any missing argument needed for function execution based on the API’s requirements, they are encouraged to ask questions to the user for this information. AwN prompts LLMs not to generate any API call until obtaining all the necessary information. In other words, only if no further information is needed, they can bypass the query step and directly initiate the API call. We also provide various kinds of specific instructions and demonstration examples for different kinds of instruction issues. All the details can be found in the Appendix due to the space limitation.
Figure 3: Illustration of Auto-Interaction module.
Experiments
In this section, we evaluate the performance of our Ask-when-Needed (AwN) prompting technique on the NoisyToolBench dataset. We first introduce the evaluation metrics, where we specify the criteria used to assess the effectiveness of AwN. Following that, we describe the evaluation pipeline, detailing the step-by-step process employed to measure AwN’s performance. Lastly, we discuss the main experiments, presenting the results and findings from our comprehensive testing of the AwN technique.
Evaluation Metrics
@@
Auto-Evaluation Pipeline
@@
Auto-Interaction ToolEvaluator can proxy the user’s communication with LLMs. When LLMs post a question, ToolEvaluator calculates the semantic similarity between the asked question and the expected question by the sentence-transformer (Reimers and Gurevych 2019). If the similarity is higher than a threshold, ToolEvaluator replies with the predefined answer to the LLMs. Otherwise, this question is treated as an irrelevant question and ToolEvaluator replies with a standard reply of ”Sorry, I cannot provide additional information about this.”. This approach streamlines the evaluation process by reducing the need for human interaction with LLMs, as illustrated in Figure 3.
Auto-Assessment ToolEvaluator can also automatically assess how well LLMs perform under ambiguous instructions according to the five metrics introduced above. A1 measures whether LLMs can ask the right question. ToolEvaluator calculates the semantic similarity between the LLMs-asked question and the expected question to asses A1. A2 measures whether LLMs can conduct correct API calls. Following the previous works (Yang et al. 2023b; Chiang and yi Lee 2023; Wang et al. 2023; Yuan et al. 2023), ToolEvaluator adopts GPT-4 as a judge to identify whether the generated API calls are the same as the expected API calls. A3 measures whether LLMs can correctly generate the final answer. ToolEvaluator adopts GPT-4 as a judge to identify whether the final answer aligns with the user’s intent. For measuring the efficiency, ToolEvaluator counts the number of generated irrelevant questions as Re and counts the total number of actions during the process as Steps. All the details can be found in the Appendix due to the space limitation.
Table 2: Assessing the accuracy of various LLMs using different prompting methods in our benchmark
base model framework IMKI IMR IwE IBTC A1(%) A2(%) A3(%) A1(%) A2(%) A3(%) A1(%) A2(%) A3(%) A1(%) A2(%) A3(%) gpt-3.5-turbo CoT 0.34 - - 0.24 - - 0.21 - - 0.00 - - + AwN 0.86 0.56 0.48 0.76 0.42 0.24 0.58 0.28 0.16 0.08 - - DFSDT 0.53 - - 0.42 - - 0.42 - - 0.00 - - + AwN 0.84 0.52 0.44 0.88 0.46 0.42 0.86 0.40 0.26 0.13 - - gpt-4-0125 CoT 0.68 - - 0.58 - - 0.35 - - 0.29 - - + AwN 0.84 0.50 0.44 0.76 0.42 0.34 0.42 0.30 0.26 0.98 - - DFSDT 0.64 - - 0.56 - - 0.37 - - 0.63 - - + AwN 0.80 0.70 0.46 0.64 0.42 0.36 0.51 0.58 0.40 0.92 - -
@@
base model framework IMKI IMR IwE IBTC Re Steps Re Steps Re Steps Re Steps gpt-3.5-turbo CoT - 4.64 - 4.52 - 5.09 - 3.10 + AwN 0.84 5.80 1.46 6.16 1.67 6.28 - 3.75 DFSDT - 50.20 - 55.50 - 45.83 - 17.06 + AwN 10.90 68.60 11.54 74.50 15.30 79.80 - 9.25 gpt-4-0125 CoT - 3.00 - 2.92 - 2.79 - 1.58 + AwN 0.18 3.70 0.20 3.40 0.23 3.47 - 1.17 DFSDT - 5.26 - 6.88 - 6.18 - 2.65 + AwN 0.34 5.30 0.36 4.98 0.37 6.76 - 1.25
The Effectiveness of ToolEvaluator
@@
Since ToolEvaluator is an automatic evaluation method, the evaluation can be inaccurate due to the imperfect nature of AI techniques, such as sentence transformer or GPT-4 as the judge. In this section, we conduct a human annotation to validate the effectiveness of ToolEvaluator. Specifically, we randomly select 50 cases and ask annotators to assess the accuracy and efficiency, according to the evaluation metrics mentioned above. Then we compare the assessment results from ToolEvaluator and human annotators. ToolEvaluator achieves 90% accuracy, indicating its effectiveness.
Experimental Setup
@@
We evaluated the performance of AwN against two baseline methods, chain-of-thought (CoT) (Wei et al. 2022) and depth-first search-based decision tree (DFSDT) (Qin et al. 2023b), which are two widely-used tool-learning methods. All the experiments are conducted with two LLMs as engines, gpt-3.5-turbo-0613 and gpt-4-0125, using the default setting. The baseline prompting methods did not integrate additional user-LLM communication during the whole execution process, leading to inevitable failures under unclear instructions. Therefore, their performance in A2 and A3 are not measured. And since an ideal reaction under Instructions Beyond Tool Capabilities (IBTC) is telling the user that the query is beyond the capabilities and refusing to generate API calls, its performance in A2 and A3 are measured neither.
Main Result
@@
We evaluate the performance of AwN as well as the baseline methods on our NoisyToolBench dataset. The accuracy-related results are shown in in Table 2 and the efficiency-related results are shown in Table 3.
@@
AwN enhances the capability of gpt-3.5-turbo to ask pertinent questions across different issues. For example, as is shown in Table 2, AwN improved the A1 scores from 0.34 to 0.86, from 0.24 to 0.76, from 0.21 to 0.58, and from 0.00 to 0.08 for gpt-3.5-turbo-based CoT as well as from 0.53 to 0.84, from 0.42 to 0.88, from 0.42 to 0.86 and from 0.00 to 0.13 for gpt-3.5-turbo-based DFSDT.
However, this improvement comes with a side effect: gpt-3.5-turbo tends to ask more irrelevant or redundant questions, as indicated by the higher Re scores in Table 3. Especially in gpt-3-turbo-based DFSDT, where the average number of redundant questions is 10.9, 11.54, and 15.3. This suggests that while the AwN aids in identifying and addressing ambiguities in user instructions, it also leads to a less efficient querying process
AwN improves gpt-4 effectively and efficiently. When gpt-4 is enhanced with the AwN method, it also demonstrates a strong capability of asking relevant questions to address ambiguity in instructions. This is evidenced by the increasing A1 scores in Table 2, from 0.68 to 0.84, from 0.58 to 0.76, from 0.35 to 0.42 and from 0.29 to 0.98 for CoT, as well as the increase from 0.64 to 0.8, from 0.56 to 0.64, from 0.37 to 0.51 and 0.63 to 0.92.
Besides, AwN can improve gpt-4 without generating excessive unnecessary questions. As is shown in Table 3, AwN only generates 0.18, 0.2, and 0.23 redundant questions for CoT, as well as 0.34, 0.36, and 0.37 redundant questions for DFSDT. Compared with gpt-3.5-turbo, gpt-4 demonstrates a stronger ability to avoid asking unnecessary questions while effectively identifying and addressing ambiguities in user instructions.
@@
Asking the right question leads to the better generation and execution of API calls. Besides the significant improvements on A1, AwN also achieves considerable performance in generating correct API calls (A2) and returning the expected final answer (A3). For example, gpt-4-based DFSDT can achieve 0.46, 0.36, and 0.40 accuracy on A3 with the help of AwN method.
@@
Impact of AwN on the average number of steps. The average number of steps measures the cost of LLMs’ tool-using. As is shown in Table 3, adopting AwN will enviably increase the number of actions. However, considering the significant performance improvements achieved, the moderate increase in cost is justifiable and worthwhile.
Case Study
Figure 4 shows a case study to illustrate the effectiveness of our AwN method. Specifically, AwN identifies that the instruction provided by the user is incorrect. Then it generates questions to communicate with users. After getting the correct information, LLM can generate correct API Call and get the expected results.
Figure 4: Case Study.
Conclusion
@@
This study delves into the critical issue of unclear user instructions and their impact on the effective use of tools by modern LLMs. Recognizing the limitations of LLMs in dealing with ambiguous instructions, we conducted a thorough investigation into the common error patterns present in real-world user instructions. Based on our analysis, we introduced the Noisy ToolBench dataset (NoisyToolBench), a novel tool-using benchmark aimed to evaluate the LLM’s tool-using performance under unclear user instructions. Furthermore, we developed the Ask-when-Needed method (AwN), an innovative approach that empowers LLMs to actively seek user input whenever they face uncertainty in instructions. To automate the evaluation process of LLMs in tool usage, we also present an automated evaluation tool (ToolEvaluator), which can proxy human users to interact with LLMs and assess LLMs’ from both accuracy and efficiency perspectives. Our experimental results show that the AwN algorithm markedly surpasses existing methods in the NoisyToolBench dataset and significantly improves the performance of LLMs’ tool-using under unclear user instructions.

### v3

@@ 4 Ask-when-Needed Prompting
AwN is built upon widely-used tool-using methods, such as CoT and ReAct. As in Figure 2, we introduce an additional step before the generation of API calls. This step involves presenting all available information to the LLMs, including the user query and API guideline, and prompting them to determine the adequacy and correctness of user instruction. If LLMs identify any missing argument needed for function execution based on the API’s requirements, they are encouraged to ask questions to the user for this information. AwN prompts LLMs not to generate any API call until obtaining all the necessary information. In other words, only if no further information is needed, they can bypass the query step and directly initiate the API call. We also provide various kinds of specific instructions and demonstration examples for different kinds of instruction issues.
@@
Figure 3: Illustration of the Auto-Interaction module.
5 Experiments
In this section, we evaluate the performance of our Ask-when-Needed (AwN) prompting technique on the NoisyToolBench dataset. We first introduce the evaluation metrics, where we specify the criteria used to assess the effectiveness of AwN. Then, we describe the evaluation pipeline, detailing the step-by-step process employed to measure AwN’s performance. Lastly, we discuss the main experiments, presenting the results and findings from our comprehensive testing of the AwN technique.
5.1 Evaluation Metrics
@@
Model Framework IMKI IMR IwE IBTC A1(%) A2(%) A3(%) A1(%) A2(%) A3(%) A1(%) A2(%) A3(%) A1(%) gpt-3.5 CoT 0.74 0.36 0.22 0.20 0.24 0.12 0.5 0.24 0.16 0.38 + AwN 0.74 0.44 0.24 0.86 0.46 0.20 0.74 0.48 0.28 0.48 DFSDT 0.64 0.16 0.12 0.60 0.18 0.16 0.48 0.14 0.14 0.46 + AwN 0.88 0.52 0.46 0.88 0.56 0.48 0.72 0.42 0.36 0.64 gpt-4 CoT 0.74 0.48 0.32 0.72 0.52 0.36 0.52 0.26 0.24 0.34 + AwN 0.94 0.62 0.50 0.76 0.44 0.38 0.48 0.34 0.34 0.94 DFSDT 0.82 0.16 0.16 0.70 0.28 0.26 0.54 0.12 0.10 0.54 + AwN 0.80 0.56 0.48 0.80 0.50 0.44 0.52 0.38 0.36 0.94 gpt-4o CoT 0.52 0.48 0.34 0.18 0.28 0.16 0.12 0.12 0.10 0.10 + AwN 0.90 0.58 0.36 0.80 0.46 0.30 0.60 0.44 0.32 0.92 DFSDT 0.58 0.20 0.18 0.26 0.18 0.16 0.18 0.06 0.04 0.08 + AwN 0.88 0.60 0.46 0.90 0.52 0.36 0.64 0.46 0.38 0.94 deepseek-v3 CoT 0.44 0.40 0.20 0.24 0.28 0.24 0.10 0.14 0.14 0.30 + AwN 0.70 0.52 0.36 0.70 0.54 0.46 0.40 0.30 0.26 0.98 DFSDT 0.42 0.30 0.26 0.60 0.20 0.18 0.22 0.12 0.12 0.48 + AwN 0.72 0.52 0.42 0.82 0.52 0.48 0.54 0.38 0.36 0.98 gemini-1.5 CoT 0.22 0.18 0.10 0.22 0.10 0.02 0.08 0.12 0.06 0.52 + AwN 0.86 0.40 0.18 0.74 0.24 0.08 0.58 0.28 0.22 0.68 DFSDT 0.62 0.02 0.02 0.6 0.08 0.04 0.36 0.06 0.02 0.48 + AwN 0.82 0.40 0.12 0.76 0.28 0.04 0.66 0.36 0.26 0.70 claude-3.5 CoT 0.24 0.26 0.20 0.12 0.28 0.24 0.08 0.26 0.24 0.30 + AwN 0.54 0.5 0.5 0.32 0.30 0.24 0.34 0.34 0.26 0.88 DFSDT 0.26 0.18 0.14 0.12 0.18 0.18 0.12 0.20 0.18 0.62 + AwN 0.52 0.44 0.42 0.32 0.30 0.18 0.36 0.36 0.30 0.86
Table 2: Assessing the accuracy of various LLMs using different prompting methods in our benchmark.
5.2 Auto-Evaluation Pipeline
@@
Auto-Interaction. ToolEvaluator can proxy the user’s communication with LLMs. When LLMs post a question, ToolEvaluator calculates the semantic similarity between the asked question and the expected question by the sentence-transformer Reimers and Gurevych (2019). If the similarity is higher than a threshold, ToolEvaluator replies with the predefined answer to the LLMs. Otherwise, this question is treated as an irrelevant question and ToolEvaluator replies with a standard reply of "Sorry, I cannot provide additional information about this.". This approach streamlines the evaluation process by reducing the need for human interaction with LLMs, as illustrated in Figure 3.
Auto-Assessment. ToolEvaluator can also automatically assess how well LLMs perform under ambiguous instructions according to the five metrics introduced above. A1 measures whether LLMs can ask the right question. ToolEvaluator calculates the semantic similarity between the LLMs-asked question and the expected question to asses A1. A2 measures whether LLMs can conduct correct API calls. Following the previous works Yang et al. (2023b); Chiang and yi Lee (2023); Wang et al. (2023); Yuan et al. (2023), ToolEvaluator adopts GPT-4o as a judge to identify whether the generated API calls are the same as the expected API calls. A3 measures whether LLMs can correctly generate the final answer. ToolEvaluator adopts GPT-4o as a judge to identify whether the final answer aligns with the user’s intent. For measuring the efficiency, ToolEvaluator counts the number of generated irrelevant questions as Re and counts the total number of actions during the process as Steps.
Model FrWork IMKI IMR IwE IBTC
Re Steps Re Steps Re Steps Re Steps
gpt-3.5 CoT - 4.46 - 4.02 - 3.90 - 2.10
+ AwN 0.66 5.36 1.1 5.98 0.76 5.08 - 2.40
DFSDT - 12.82 - 12.80 - 13.82 - 5.50
+ AwN 1.44 16.94 0.98 11.24 0.94 11.68 - 3.94
gpt-4 CoT - 4.00 - 3.98 - 3.34 - 2.04
+ AwN 0.16 3.94 0.20 3.94 0.36 3.46 - 1.16
DFSDT - 83.96 - 21.04 - 22.40 - 4.06
+ AwN 0.48 9.82 0.74 13.08 0.62 9.42 - 2.10
gpt-4o CoT - 3.00 - 2.98 - 2.48 - 1.28
+ AwN 0.62 3.86 0.70 3.96 0.46 3.18 - 1.10
DFSDT - 5.98 - 9.58 - 5.78 - 8.98
+ AwN 0.86 6.70 1.18 7.68 0.88 8.56 - 1.14
deepseek-v3 CoT - 4.20 - 3.52 - 3.12 - 1.18
+ AwN 0.20 3.88 0.06 3.60 0.04 2.92 - 1.10
DFSDT - 59.08 - 41.70 - 24.24 - 11.64
+ AwN 1.16 15.86 1.80 24.60 1.20 11.82 - 1.32
gemini-1.5 CoT - 4.00 - 4.12 - 2.86 - 4.52
+ AwN 0.42 6.44 0.68 6.36 0.48 4.54 - 1.46
DFSDT - 750.80 - 685.00 - 725.14 - 559.78
+ AwN 5.34 445.16 9.08 532.56 1.94 411.18 - 1.46
claude-3.5 CoT - 2.64 - 3.40 - 3.04 - 1.90
+ AwN 0.18 3.74 0.33 1.03 0.36 3.76 - 1.68
DFSDT - 3.34 - 9.64 - 5.98 - 4.26
+ AwN 0.76 6.74 0.80 17.46 1.04 13.08 - 2.76
@@
5.3 The Effectiveness of ToolEvaluator
@@
Since ToolEvaluator is an automatic evaluation method, the evaluation can be inaccurate due to the imperfect nature of AI techniques, such as sentence transformer or GPT-4o as the judge. In this section, we conduct a human annotation to validate the effectiveness of ToolEvaluator. Specifically, we randomly select 50 cases and ask annotators to assess the accuracy and efficiency, according to the evaluation metrics mentioned above. Then we compare the assessment results from ToolEvaluator and human annotators. ToolEvaluator achieves 90% accuracy, indicating its effectiveness.
5.4 Experimental Setup
@@
We evaluated the performance of AwN against two baseline methods, chain-of-thought (CoT) Wei et al. (2022) and depth-first search-based decision tree (DFSDT) Qin et al. (2023b), which are two widely-used tool-learning methods. All the experiments are conducted with several LLMs as engines, gpt-3.5-turbo-0125, gpt-4-turbo-2024-04-09, gpt-4o-2024-11-20, deepseek-v3, gemini-1.5-flash-latest and claude-3-5-haiku-20241022, using the default setting. Since an ideal reaction under Instructions Beyond Tool Capabilities (IBTC) is telling the user that the query is beyond the capabilities and refusing to generate API calls, its performance in A2 and A3 are measured neither.
5.5 Main Result
@@
We evaluate the performance of AwN as well as the baseline methods on our NoisyToolBench dataset. The accuracy-related results are shown in Table 2 and the efficiency-related results are in Table 3.
@@
AwN enhances the capability of LLM Agents to ask pertinent questions across different issues. For example, as is shown in Table 2, AwN improved the A1 scores from 0.52 to 0.90, from 0.18 to 0.80, from 0.12 to 0.60, and from 0.10 to 0.92 for gpt-4o-based CoT as well as from 0.58 to 0.88, from 0.26 to 0.90, from 0.18 to 0.64 and from 0.08 to 0.94 for gpt-4o-based DFSDT.
@@
Asking the right question leads to the better generation and execution of API calls. Besides the significant improvements on A1, AwN also achieves considerable performance in generating correct API calls (A2) and returning the expected final answer (A3). For example, AwN improved the A2 scores from 0.48 to 0.58, from 0.28 to 0.46, from 0.12 to 0.44 for gpt-4o-based CoT as well as from 0.20 to 0.60, from 0.18 to 0.52, from 0.06 to 0.46 for gpt-4o-based DFSDT.
AwN can improve most of the LLM agents without generating excessive unnecessary questions. As is shown in Table 3, AwN only leads to 0.16, 0.20, and 0.36 redundant questions for gpt-4-based-CoT, as well as 0.48, 0.74, and 0.62 redundant questions for gpt-4-based-DFSDT.
However, a few LLM agents tend to ask more irrelevant or redundant questions, as indicated by the higher Re scores in Table 3. For example, in Gemini-1.5-based DFSDT, where the average number of redundant questions is 5.34, 9.08, and 1.94. This suggests that while the AwN aids in identifying and addressing ambiguities in user instructions, it also leads to a less efficient querying process.
@@
AwN can reduce the average cost of LLM’s tool-using. The average number of steps measures the cost of LLMs’ tool-using. As is shown in Table 3, adopting AwN can reduce the number of actions, especially for gpt-4-based DFSDT, deepseek-v3-based-DFSDT and gemini-1.5-based-DFSDT. Although AwN can lead to a higher cost for a few LLM agents, such as claude-3.5, considering the significant performance improvements achieved, the moderate increase in cost is justifiable and worthwhile.
6 Conclusion
@@
This study explores how unclear user instructions hinder modern LLMs’ tool usage. To investigating the common error patterns in real-world instructions, we propose: (1) Noisy ToolBench (NoisyToolBench), a novel benchmark for evaluating LLM performance under ambiguous instructions; (2) Ask-when-Needed (AwN), an innovative approach enabling LLMs to request clarification when uncertain; and (3) an automated evaluator (ToolEvaluator) to assess accuracy and efficiency. Experimental results show that the AwN algorithm markedly surpasses existing methods in the NoisyToolBench dataset and significantly improves the performance of LLMs’ tool-using under unclear user instructions.